# Discovering new robust local search algorithms with neuro-evolution

### v1

@@ Acknowledgment
Conflict of Interest
On behalf of all authors, the corresponding author, Olivier Goudet, states that there is no conflict of interest.
@@
Authors’ contributions
MS. Amri Sakhri, O. Goudet and C. Touhami developed the code, and performed the experiments. A. Goëffon and F. Saubion planned and supervised the work. All the authors contributed to the analysis and the writing of the manuscript.
Data Availability Statement
The instances used in this work and the source code of the tested methods are available at the url https://github.com/Salim-AMRI/NK_Landscape_Project.git.
Research Involving Human and /or Animals
Not Applicable
Informed Consent
Not Applicable
@@ Appendix A Other emerging strategies
In this section we analyze the other strategies that emerged when the $o_{2}$ and $o_{4}$ observation functions were used. Analyses of these strategies are a little less easy to interpret, as in this case the neural network learned is a non-linear function of two input variables instead of one.
@@
Figure 9 displays a representative example of the trajectory performed by Neuro-LS using observation $o_{2}$ when $N=64$ and $K=8$. We observe on this plot that the emerging Neuro-LS strategy has two successive operating modes like the strategy based on the variations of fitness for each possible move (see Section 4.5.2): small steps hill climbing behavior and Jump with worst move. It is interesting to see that this same behavior can also emerge when using very raw information as input.
@@
Figure 9: Fitness evolution curve and strategy used by Neuro-LS using $o_{2}$ observations for the resolution of a instance with significantly rugged NK landscape ($N=64$ and $K=8$).
We display in Figure 10 the continuous function $g_{\theta}$ learned by the neural network taken a 2-dimensional vector as input. We see in this figure that the Neuro-LS strategy also gives the worst preference in this case to moves that deteriorate slightly, when the new fitness (after the move) is slightly lower than the current fitness (red dots on the graph). At the top of the graph, we see in green that the most preferred moves are the small improvement moves, or strong deterioration moves, when the new fitness after the flip is really inferior to the current fitness (allowing for jumping moves).
Figure 10: Output of the neural network $g_{\theta}(o_{2}(x)_{i}))$ (z-axis) for each 2-dimensional input with the current fitness $f_{{\rm obj}}(x)$ (x-axis) and the new fitness $f_{{\rm obj}}(flip_{i}(x))$ (y-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$.
@@
Figure 11 displays a representative example of the trajectory performed by Neuro-LS using observation $o_{4}$ when $N=64$ and $K=8$. We observe on this plot that the emerging Neuro-LS strategy has also two successive operating modes like the other strategies: small steps hill climbing behavior and jump with worst move.
@@
On the left of this figure, we can also see that very deteriorant moves with a very negative z-score also have a fairly favorable score, as they are the ones that allow Neuro-LS to make a jump when there are no more improving moves.
Figure 11: Fitness evolution curve and strategy used by Neuro-LS using $o_{4}$ observations for the resolution of a instance with significantly rugged NK landscape ($N=64$ and $K=8$).
Figure 12: Output of the neural network $g_{\theta}(o_{4}(x)_{i}))$ (z-axis) for each 2-dimensional input with the ranking score $o_{3}(x)_{i}$ (x-axis) and the z-score $Z(\Delta_{i}(x))$ (y-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$.

### v2

@@ Appendix A Other emerging strategies
In this section we analyze the other strategies that emerged when the $o^{2}$ and $o^{4}$ observation functions were used. Analyses of these strategies are a little less easy to interpret, as in this case the neural network learned is a non-linear function of two input variables instead of one.
@@
Figure 9 displays a representative example of the trajectory performed by Neuro-LS using observation $o^{2}$ when $N=64$ and $K=8$. We observe on this plot that the emerging Neuro-LS strategy has two successive operating modes like the strategy based on the variations of fitness for each possible move (see Section 4.5.2): small steps hill climbing behavior and Jump with worst move. It is interesting to see that this same behavior can also emerge when using very raw information as input.
@@
Figure 9: Fitness evolution curve and strategy used by Neuro-LS using $o^{2}$ observations for the resolution of a instance with significantly rugged NK landscape ($N=64$ and $K=8$).
We display in Figure 10 the continuous function $g_{\theta}$ learned by the neural network taken a 2-dimensional vector as input. We see in this figure that the Neuro-LS strategy also gives the worst preference in this case to moves that deteriorate slightly, when the new fitness (after the move) is slightly lower than the current fitness (red hexagons on the heatmap). We see in dark green that the most preferred moves are the small improvement moves. The strong deterioration moves have also a quite good preference score, when the new fitness after the flip is really inferior to the current fitness (allowing for jumping moves).
Figure 10: Output of the neural network $g_{\theta}(o^{2}(x)_{i})$, or preference score for the move $i$, for each 2-dimensional input with the current fitness $f_{{\rm obj}}(x)$ (x-axis) and the new fitness $f_{{\rm obj}}(flip_{i}(x))$ (y-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$.
@@
Figure 11 displays a representative example of the trajectory performed by Neuro-LS using observation $o^{4}$ when $N=64$ and $K=8$. We observe on this plot that the emerging Neuro-LS strategy has also two successive operating modes like the other strategies: small steps hill climbing behavior and jump with worst move.
@@
On the left of this figure, we can also see that very deteriorating moves with a very negative z-score also have a fairly favorable score, as they are the ones that allow Neuro-LS to make a jump when there are no more improving moves.
Figure 11: Fitness evolution curve and strategy used by Neuro-LS using $o^{4}$ observations for the resolution of a instance with significantly rugged NK landscape ($N=64$ and $K=8$).
Figure 12: Output of the neural network $g_{\theta}(o^{4}(x)_{i})$, or preference score for the move $i$, for each 2-dimensional input with the ranking score $o^{3}(x)_{i}$ (x-axis) and the z-score $Z(\Delta_{i}(x))$ (y-axis). It corresponds to the values collected during 10 trajectories of Neuro-LS on the test instances of the NK Landscape problem with $N=64$ and $K=8$.